rxcklesskaisher:

Which statement best describes the area of the triangle shown below? A coordinate grid is shown with a triangle.. The base is 4 units, and the height is 4 units. It is twice the area of a rectangle of length 4 units and width 2 units. It is one-half the area of a square of side length 4 units. It is one-half the area of a rectangle of length 4 units and width 2 units. It is twice the area of a square of side length 4 units.

supie:

A good way to do this would be to find the area of the triangle, it has already given that the base is 4 and the height is 4 we use \(a=\frac{h_bb}{2}\) to fidn the area of a triangle h=height and b= base the base is 4 and the height is 4 so plugging in the values into the formula we have \(\frac{4 \times 4}{2}=?\) so solve that by multiplying 4 times 4 then dividing that by 2